Circular No. 8037

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002jz IN UGC 2984 T. Puckett, Mountain Town, GA; and D. Tigner, Kanata, ON, report

the discovery of an apparent supernova (mag 15.9) on an unfiltered CCD frame (limiting mag 18.8) taken with the Puckett Observatory 0.35-m automated supernova patrol telescope on Dec. 23.27 UT. The new object is located at α = 4^{h}13^{m}12^{s}.52, δ = +13^{o}25′07′′.3 (equinox 2000.0), which is 1′′.9 west and 4′′.2 south of the center of UGC 2984. SN 2002jz was confirmed on unfiltered CCD frames taken by A. Sehgal, Woodinville, WA, on Dec. 24.27, but it does not appear on unfiltered images taken by Puckett on 2001 Feb. 26, Sept. 22, and Dec. 5 (limiting mag ∼ 19.5), or on Palomar Sky Survey images taken on 1990 Jan. 24, 1989 Nov. 5 (limiting mag ∼ 21.0), and 1953 Oct. 10 (limiting mag ∼ 20.0).

M. Hamuy, Carnegie Observatories, reports spectroscopic observations (range 380–725 nm) of SN 2002jz made by M. Roth and N. Morrell on Dec. 27.22 UT with the Las Campanas Observatory Dupont 2.5-m telescope (+ Modular Spectrograph). The spectrum bears some resemblance to that of the type-Ic SN 1994I taken 10 days past maximum (Filippenko et al. 1995, Ap.J. 450, L11). Both supernovae show a strong Na I D 589.2-nm line exhibiting a P-Cyg profile and prominent blends of Fe II lines between 400 and 470 nm. However, SN 2002jz shows an evident absorption (possibly Hα) at 644 nm (rest wavelength) that is not present in SN 1994I, suggesting perhaps an intermediate IIb-type supernova. Na I D interstellar lines at the redshift of the host galaxy can be seen with an equivalent width of 0.1 nm, which suggests significant reddening by dust.

SUPERNOVA 2002jy IN NGC 477 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002jy (cf. IAUC 8035), obtained by M. Calkins on Dec. 26.11 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 5876 km/s for the host galaxy, the supernova expansion velocity is ∼ 11000 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days before maximum light.

2002 December 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green